Solve the following system of equations.
x+3y=3 (1)
x=4-3y (2)
What is the solution of the system?
A. The solution is______? (type it as an ordered pair.)
B. No solution.

Answer by htmentor(1343)

Solve the following system of equations.
x+3y=3 (1)
x=4-3y (2)
What is the solution of the system?
Adding 2y to both sides of equation (2) gives
x+3y=4
It is clear from this that there is no solution, since we have
x+3y=3
x+3y=4